# Rank-1 Approximation of Inverse Fisher for Natural Policy Gradients in Deep Reinforcement Learning

### v1

@@ Appendix A Appendix
A.1 Assumptions
@@
$\bm{F}_{\rho}(\theta)=\operatorname{\mathbb{E}}_{s\sim d_{\rho}^{\pi_{\theta}}}\operatorname{\mathbb{E}}_{a\sim\operatorname{\pi_{\theta}}}\left[\nabla_{\theta}\log\pi_{\theta}\cdot\nabla_{\theta}\log\pi_{\theta}^{T}\right]\curlyeqsucc\mu_{F}\mathbf{I}$ (8)
@@
where $\bm{w}^{\theta}_{\star}=\operatorname*{argmin}_{w}L_{\nu^{\pi_{\theta}}_{\rho}}(\bm{w};\theta)$ is obtained via the full natural policy gradient direction at the parametrization $\theta$.
@@
A.2 Helper Lemmas
@@
Liu et al. (2020) In the stochastic PG update, by choosing $\eta=\frac{1}{4L_{J}}$, $K=\frac{32L_{J}(J^{H,\star}-J^{H}(\theta_{0}))}{\varepsilon},$ and $N=\frac{6\sigma^{2}}{\varepsilon}$, we have
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla J^{H}(\theta^{k})\|^{2}]\leq\varepsilon.$
@@
Xu et al. (2020) Let $J(\theta),J^{H}(\theta)$ are $L_{J}$-smooth, where $L_{J}=\frac{MR}{(1-\gamma)^{2}}+\frac{2G^{2}R}{(1-\gamma)^{3}}$ and
@@
Liu et al. (2020) Let $g^{k}=\frac{1}{N}\sum_{i=1}^{N}g(\tau^{H}_{i}|\theta^{k})$ be the policy gradient estimate for samples until horizon H. Then, we have
Lemma 4.
Liu et al. (2020) Let $\bm{w}^{k}_{\star}=F_{\rho}^{-1}(\theta^{k})\nabla J(\theta^{k})$ be the exact NPG update direction at $\theta^{k}$. Then, we have
@@
A.3 Global Convergence of SM Update
@@
    We know from Jensen’s inequality and $\left(\mathbb{E}[\|\bm{w}^{j+1}_{t}-\bm{w}^{j+1}_{t,\star}\|]\right)^{2}\leq\mathbb{E}[\|\bm{w}^{j+1}_{t}-\bm{w}^{j+1}_{t,\star}\|^{2}]$ that
    $\begin{split}&\left(\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\bm{w}^{k}-\bm{w}^{k}_{\star}\|]\right)^{2}\\ &\leq\frac{1}{K}\sum_{k=0}^{K-1}\left(\mathbb{E}[\|\bm{w}^{k}-\bm{w}^{k}_{\star}\|]\right)^{2}\\ &\leq\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\bm{w}^{k}-\bm{w}^{k}_{\star}\|^{2}]\\ &\leq\frac{2}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\bm{w}^{k}-\nabla J(\theta^{k})\|^{2}]+\frac{2}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla J(\theta^{k})-\bm{w}^{k}_{\star}]\end{split}$ (14)
@@
    be the weight update from the Sherman-Morrison update then we have from Lemma 3 and Assumption 2 that
    $\begin{split}&\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\bm{w}^{k}-\nabla J(\theta^{k})\|^{2}]\\ &=\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\left\lVert\frac{1}{\lambda}\bm{g}^{k}-\frac{\bm{\hat{F}}^{k}}{\lambda^{2}+\lambda Tr(\hat{\bm{F}}^{k})}\bm{g}^{k}-\nabla J(\theta^{k})\right\lVert^{2}\right]\\ &=\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\left\lVert\frac{1}{\lambda}(\bm{g}^{k}-\nabla J^{H}(\theta^{k}))-\frac{\bm{\hat{F}}^{k}}{\lambda^{2}+\lambda Tr(\hat{\bm{F}}^{k})}(\bm{g}^{k}-\nabla J^{H}(\theta^{k}))-(\nabla J(\theta^{k})-\nabla J^{H}(\theta^{k}))\right.\right.\\ &\,\,\,\left.\left.+\frac{1}{\lambda}\nabla J^{H}(\theta^{k})-\frac{\bm{\hat{F}}^{k}}{\lambda^{2}+\lambda Tr(\hat{\bm{F}}^{k})}\nabla J^{H}(\theta^{k})-\nabla J^{H}(\theta^{k})\right\lVert^{2}\right]\\ &\leq 2\frac{\sigma^{2}}{N}\left(\frac{1}{\lambda^{2}}+\frac{G^{2}}{\lambda^{2}+\lambda\mu_{F}}\right)+2G^{2}R^{2}\left(\frac{H+1}{1-\gamma}+\frac{\gamma}{(1-\gamma)^{2}}\right)^{2}\gamma^{2H}\\ &\,\,\,+2\left(1+\frac{1}{\lambda^{2}}+\frac{G^{2}}{\lambda^{2}+\lambda\mu_{F}}\right)\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla J^{H}(\theta^{k})\|^{2}]\end{split}$ (15)
@@
    $\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla J^{H}(\theta^{k})\|^{2}]\leq\frac{J^{H\star}-J^{H}(\theta_{0})}{K}+\frac{3\sigma^{2}}{N}$
@@
    $\begin{split}&\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\|\bm{w}^{k}\|^{2}\\ &=\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\left\lVert\frac{1}{\lambda}\bm{g}^{k}-\frac{\bm{\hat{F}}^{k}}{\lambda^{2}+\lambda Tr(\hat{\bm{F}}^{k})}\bm{g}^{k}\right\lVert^{2}\right]\\ &=\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\left\lVert\frac{1}{\lambda}(\bm{g}^{k}-\nabla J^{H}(\theta^{k}))-\frac{\bm{\hat{F}}^{k}}{\lambda^{2}+\lambda Tr(\hat{\bm{F}}^{k})}(\bm{g}^{k}-\nabla J^{H}(\theta^{k}))+\frac{1}{\lambda}\nabla J^{H}(\theta^{k})\right.\right.\\ &\,\,\,\left.\left.-\frac{\bm{\hat{F}}^{k}}{\lambda^{2}+\lambda Tr(\hat{\bm{F}}^{k})}\nabla J^{H}(\theta^{k})\right\lVert^{2}\right]\\ &\leq 2\frac{\sigma^{2}}{N}\left(\frac{1}{\lambda^{2}}+\frac{G^{2}}{\lambda^{2}+\lambda\mu_{F}}\right)+2\left(\frac{1}{\lambda^{2}}+\frac{G^{2}}{\lambda^{2}+\lambda\mu_{F}}\right)\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla J^{H}(\theta^{k})\|^{2}]\\ &\leq 8\frac{\sigma^{2}}{N}\left(\frac{1}{\lambda^{2}}+\frac{G^{2}}{\lambda^{2}+\lambda\mu_{F}}\right)+\frac{16L_{J}J^{H\star}-J^{H}(\theta_{0})}{K}\left(\frac{1}{\lambda^{2}}+\frac{G^{2}}{\lambda^{2}+\lambda\mu_{F}}\right).\end{split}$
@@
    $\begin{split}N&\geq\frac{8\eta\sigma^{2}\left(\frac{1}{\lambda^{2}}+\frac{G^{2}}{\lambda^{2}+\lambda\mu_{F}}\right)}{\frac{\varepsilon}{6}},\\ K&\geq\frac{16L_{J}\eta(J^{H,\star}-J^{H}(\theta_{0}))\left(\frac{1}{\lambda^{2}}+\frac{G^{2}}{\lambda^{2}+\lambda\mu_{F}}\right)}{\frac{\varepsilon}{6}},\end{split}$ (20)
@@
In summary, we require NN and $K$ to satisfy equation 18, equation 20, and equation 22, which leads to
@@
A.4 Additional Abalation Studies
Figure 4: Average log-probabilities of the actions taken on three classic control tasks. All results are averaged over 55 random seeds, with shaded areas showing standard deviation.
Figure 5: Average log-probabilities of the actions taken on six MuJoCo tasks. All results are averaged over 55 random seeds, with shaded areas showing standard deviation.
A.5 Hyper-parameters for Experiments
  • $\eta$ (actor) : step size for the policyŋ‑network update
  • $\alpha$ (critic) : learning rate for the value network, optimised with Adam

### v2

@@ 7 Conclusion
Broader Impact
Reinforcement learning is now used in many domains, from fine-tuning language models to controlling systems such as power grids or data centres, and this broad use can bring both benefits and risks. It can improve efficiency, safety, and energy use, but it can also amplify existing biases or support decision systems that have harmful social or economic effects. Our work focuses only on the optimisation aspect: we study how to solve a given policy optimisation problem faster, so that a fixed objective is reached in fewer gradient steps. If the training data for language models or the environment model for control tasks reflects real-world dynamics and values accurately, a more efficient optimiser can produce better policies with less computation and a smaller training carbon footprint. At the same time, any negative societal impact would mainly arise from biased, incomplete, or misaligned data, objectives, and reward designs, rather than from the particular optimization method used.
Acknowledgment
SPD was funded by the Dean’s Doctoral Scholarship at the University of Manchester. SK acknowledges funding from the UKRI Turing AI World-Leading Researcher Fellowship (EP/W002973/1). MS acknowledges funding from AI Hub in Generative Models by Engineering & Physical Sciences Research Council (EPSRC) with Funding Body Ref: EP/Y028805/1.
Code availability
@@ Appendix A Appendix
A.1 Notations.
Let $\gamma\in(0,1)$ be the discount, $r:\mathcal{S}\times\mathcal{A}\rightarrow[-R,R]$ is the reward function, and $H\in\mathbb{N}\cup\{\infty\}$ the horizon. Define the truncated objective
$J^{H}(\theta)\;=\;\operatorname{\mathbb{E}}_{\tau\sim\pi_{\theta}}\!\Big[\sum_{t=0}^{H-1}\gamma^{t}\,r(s_{t},a_{t})\Big],\qquad J(\theta)=J^{\infty}(\theta).$ (16)
For each trajectory $\tau^{H}=(s_{0},a_{0},\ldots,s_{H-1},a_{H-1})$ sampled under $\pi_{\theta}$, let $g(\tau^{H}\mid\theta)$ denote a single-trajectory unbiased estimator of the truncated policy gradient:
$\operatorname{\mathbb{E}}\!\left[g(\tau^{H}\mid\theta)\right]=\nabla J^{H}(\theta).$ (17)
$g(\tau^{H}\mid\theta)=\sum_{t=0}^{H-1}\gamma^{t}\,G_{t}\,\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})$ (18)
A.2 Assumptions
Assumption 1.
Let $g(\tau^{H}|\theta)$ be an unbiased single-trajectory estimator of $\nabla J^{H}(\theta)$ Assume there exists $\sigma>0$ such that for all $\theta$,
$\mathbb{E}\!\left[\left\|g(\tau^{H}\mid\theta)-\mathbb{E}\!\left[g(\tau^{H}\mid\theta)\right]\right\|^{2}\right]\leq\sigma^{2},$
@@ A.2 Assumptions
$\bm{F}(\theta)=\operatorname{\mathbb{E}}_{s\sim d_{\rho}^{\pi_{\theta}}}\operatorname{\mathbb{E}}_{a\sim\operatorname{\pi_{\theta}}}\left[\nabla_{\theta}\log\pi_{\theta}\cdot\nabla_{\theta}\log\pi_{\theta}^{T}\right]\curlyeqsucc\mu_{F}\mathbf{I}$ (19)
@@
where $\nu^{\star}(s,a)=d_{\rho}^{\pi^{\star}}(s)\cdot\pi^{\star}(s,a)$ is the state-action distribution induced by an optimal policy $\pi^{\star}$ and $\bm{w}^{\theta}_{\star}=\operatorname*{argmin}_{w}L_{\nu^{\pi_{\theta}}_{\rho}}(\bm{w};\theta)$ is obtained via the full natural policy gradient direction at the parametrization $\theta$.
@@
Remark 3.
Assumption 1 treats $\sigma^{2}$ as a constant upper bound on the variance of the truncated policy gradient estimator, although in general $\mathrm{Var}(g(\tau^{H}\mid\theta))$ may depend on the horizon $H$. More precisely, for each finite $H$ we write
$\mathrm{Var}\big(g(\tau^{H}\mid\theta)\big)=\sigma_{H}^{2}.$
For likelihood–ratio estimators such as REINFORCE, existing analyses show that $\sigma_{H}^{2}$ can grow with $H$, but remains finite under smooth Gaussian policies with bounded rewards and features (Zhao et al., 2012; Pirotta et al., 2013; Baxter and Bartlett, 2001). In our proof we use a finite truncated horizon and define
$\sigma^{2}\coloneqq\sup_{1\leq H\leq H_{\max}}\sigma_{H}^{2}<\infty.$
This form of bounded–variance assumption is standard in recent convergence analyses of policy gradient methods (Liu et al., 2022; Xu et al., 2020; 2019).
A.3 Helper Lemmas
@@
(Liu et al., 2022) In the stochastic PG update, we have
@@
(Liu et al., 2022) Let $J(\theta),J^{H}(\theta)$ are $L_{J}$-smooth, where $L_{J}=\frac{MR}{(1-\gamma)^{2}}+\frac{2G^{2}R}{(1-\gamma)^{3}}$ and
@@
(Liu et al., 2022) Let $\bm{w}^{k}_{\star}=F^{-1}(\theta^{k})\nabla J(\theta^{k})$ be the exact NPG update direction at $\theta^{k}$. Then, we have
@@
A.4 Global Convergence of SM Update
@@
    We know from Jensen’s inequality that $\left(\mathbb{E}[\|\bm{w}^{j+1}_{t}-\bm{w}^{j+1}_{t,\star}\|]\right)^{2}\leq\mathbb{E}[\|\bm{w}^{j+1}_{t}-\bm{w}^{j+1}_{t,\star}\|^{2}]$ so,
    $\begin{split}&\left(\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\bm{w}^{k}-\bm{w}^{k}_{\star}\|]\right)^{2}\\ &\leq\frac{1}{K}\sum_{k=0}^{K-1}\left(\mathbb{E}[\|\bm{w}^{k}-\bm{w}^{k}_{\star}\|]\right)^{2}\\ &\leq\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\bm{w}^{k}-\bm{w}^{k}_{\star}\|^{2}]\\ &\leq\frac{2}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\bm{w}^{k}-\nabla J(\theta^{k})\|^{2}]+\frac{2}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla J(\theta^{k})-\bm{w}^{k}_{\star}\|^{2}]\end{split}$ (25)
@@
    be the weight update from the Sherman-Morrison update then we have from Lemma 2 and Remark 2 and Assumption 3 that
    $\begin{split}&\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\bm{w}^{k}-\nabla J(\theta^{k})\|^{2}]\\ &=\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\left\lVert\frac{1}{\lambda}\bm{g}^{k}-\frac{\bm{\hat{F}}^{k}}{\lambda^{2}+\lambda Tr(\hat{\bm{F}}^{k})}\bm{g}^{k}-\nabla J(\theta^{k})\right\lVert^{2}\right]\\ &=\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\left\lVert\frac{1}{\lambda}(\bm{g}^{k}-\nabla J^{H}(\theta^{k}))-\frac{\bm{\hat{F}}^{k}}{\lambda^{2}+\lambda Tr(\hat{\bm{F}}^{k})}(\bm{g}^{k}-\nabla J^{H}(\theta^{k}))-(\nabla J(\theta^{k})-\nabla J^{H}(\theta^{k}))\right.\right.\\ &\,\,\,\left.\left.+\frac{1}{\lambda}\nabla J^{H}(\theta^{k})-\frac{\bm{\hat{F}}^{k}}{\lambda^{2}+\lambda Tr(\hat{\bm{F}}^{k})}\nabla J^{H}(\theta^{k})-\nabla J^{H}(\theta^{k})\right\lVert^{2}\right]\\ &\leq 4\frac{\sigma^{2}}{N}\left(\frac{1}{\lambda^{2}}+\frac{G^{2}}{\lambda^{2}+\lambda\mu_{F}}\right)+4G^{2}R^{2}\left(\frac{H+1}{1-\gamma}+\frac{\gamma}{(1-\gamma)^{2}}\right)^{2}\gamma^{2H}\\ &\,\,\,+4\left(1+\frac{1}{\lambda^{2}}+\frac{G^{2}}{\lambda^{2}+\lambda\mu_{F}}\right)\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla J^{H}(\theta^{k})\|^{2}]\end{split}$ (26)
@@
    $\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla J^{H}(\theta^{k})\|^{2}]\leq\frac{16L_{J}(J^{H,\star}-J^{H}(\theta_{0}))}{K}+\frac{3\sigma^{2}}{N}$
@@
    $\begin{split}&\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\|\bm{w}^{k}\|^{2}\\ &=\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\left\lVert\frac{1}{\lambda}\bm{g}^{k}-\frac{\bm{\hat{F}}^{k}}{\lambda^{2}+\lambda Tr(\hat{\bm{F}}^{k})}\bm{g}^{k}\right\lVert^{2}\right]\\ &=\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\left\lVert\frac{1}{\lambda}(\bm{g}^{k}-\nabla J^{H}(\theta^{k}))-\frac{\bm{\hat{F}}^{k}}{\lambda^{2}+\lambda Tr(\hat{\bm{F}}^{k})}(\bm{g}^{k}-\nabla J^{H}(\theta^{k}))+\frac{1}{\lambda}\nabla J^{H}(\theta^{k})\right.\right.\\ &\,\,\,\left.\left.-\frac{\bm{\hat{F}}^{k}}{\lambda^{2}+\lambda Tr(\hat{\bm{F}}^{k})}\nabla J^{H}(\theta^{k})\right\lVert^{2}\right]\\ &\leq 2\frac{\sigma^{2}}{N}\left(\frac{1}{\lambda^{2}}+\frac{G^{2}}{\lambda^{2}+\lambda\mu_{F}}\right)+2\left(\frac{1}{\lambda^{2}}+\frac{G^{2}}{\lambda^{2}+\lambda\mu_{F}}\right)\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla J^{H}(\theta^{k})\|^{2}]\\ &\leq 16\frac{\sigma^{2}}{N}\left(\frac{1}{\lambda^{2}}+\frac{G^{2}}{\lambda^{2}+\lambda\mu_{F}}\right)+\frac{32L_{J}J^{H\star}-J^{H}(\theta_{0})}{K}\left(\frac{1}{\lambda^{2}}+\frac{G^{2}}{\lambda^{2}+\lambda\mu_{F}}\right).\end{split}$
@@
    $\begin{split}N&\geq\frac{16\eta\sigma^{2}\left(\frac{1}{\lambda^{2}}+\frac{G^{2}}{\lambda^{2}+\lambda\mu_{F}}\right)}{\frac{\varepsilon}{6}},\\ K&\geq\frac{32L_{J}\eta(J^{H,\star}-J^{H}(\theta_{0}))\left(\frac{1}{\lambda^{2}}+\frac{G^{2}}{\lambda^{2}+\lambda\mu_{F}}\right)}{\frac{\varepsilon}{6}},\end{split}$ (31)
@@
From equation 29
$\left(\frac{H+1}{1-\gamma}+\frac{\gamma}{(1-\gamma)^{2}}\right)^{2}\gamma^{2H}\;\leq\;C_{0}\,\varepsilon^{2},$ (35)
for some constant $C_{0}>0$ that does not depend on $H$ or $\varepsilon$.
For all $H\geq 0$
$\frac{H+1}{1-\gamma}+\frac{\gamma}{(1-\gamma)^{2}}\;\leq\;\frac{H+1}{1-\gamma}+\frac{1}{(1-\gamma)^{2}}\;\leq\;\frac{C_{1}(H+1)}{(1-\gamma)^{2}},$
for a constant $C_{1}\geq 1$ that depends only on $\gamma$. Substituting this into Equation 35 gives
$(H+1)^{2}\gamma^{2H}\;\leq\;C_{2}(1-\gamma)^{4}\,\varepsilon^{2},$ (36)
for constant $C_{2}>0$ independent of $H$ and $\varepsilon$.
Since the prefactor grows polynomially in $H$ while $\gamma^{2H}$ decays geometrically,
$\gamma^{2H}\leq c_{1}\,\varepsilon^{2},$
for some constant $c_{1}>0$ independent of $H$. Taking logarithm on both side yields
$H=\mathcal{O}\!\left(\log\!\big((1-\gamma)^{-1}\varepsilon^{-1}\big)\right).$ (37)
The lower bounds on NN in equation 29 and equation 31 have the form
$N\;\geq\;c_{2}\sigma^{2}\left(1+(1+\tfrac{1}{\mu_{F}})^{2}+\tfrac{1}{\lambda^{2}}+\tfrac{G^{2}}{\lambda^{2}+\lambda\mu_{F}}\right)\varepsilon^{-2},$
for a constant $c_{2}$ that absorbs numerical factors. Since the expression in parentheses is independent of $\varepsilon$, we have
$N=\mathcal{O}\!\left(\frac{\sigma^{2}}{\varepsilon^{2}}\right).$ (38)
The lower bounds on $K$ collected from equation 29, equation 31, and equation 33 can be written as
$K\;\geq\;c_{3}\left(L_{J}(J^{H,\star}-J^{H}(\theta_{0}))+\mathbb{E}_{s\sim d^{\pi^{\star}}_{\rho}}\left[\text{KL}\left(\pi^{\star}(\cdot|s)||\pi_{\theta^{0}}(\cdot|s)\right)\right]\right)\varepsilon^{-2},$
for a constant $c_{3}$ depending only on $(1-\gamma)^{-1}$ and the model parameters. Since both terms in parentheses scale as $\mathcal{O}((1-\gamma)^{-2})$. We have
$K=\mathcal{O}\!\left(\frac{1}{(1-\gamma)^{2}\varepsilon^{2}}\right).$ (39)
In summary, we require $N$ , $K$ and $H$ to satisfy equation 29, equation 31, and equation 33, which leads to
@@
A.5 Detailed Derivation of the Toy Problem
The policy is parameterized by a vector $\bm{\theta}=[\theta_{0},\theta_{1}]^{\top}$, with independent Bernoulli distribution:
$\pi_{\bm{\theta}}(a=0\mid x=0)$ $=\frac{1}{1+e^{\lambda_{p}\theta_{0}}},$ (40)
$\pi_{\bm{\theta}}(a=0\mid x=1)$ $=\frac{1}{1+e^{\theta_{1}}},$ (41)
where $\lambda_{p}$ scales the first policy parameter and is used to test scaling invariance. If we assume that the agent starts at the state $x=0$ then the objective (expected return) in this simplified setting is
$J(\bm{\theta})$ $=\pi_{\bm{\theta}}(a=0\mid x=0)+2(1-\pi_{\bm{\theta}}(a=0\mid x=0))\,\pi_{\bm{\theta}}(a=0\mid x=1).$ (42)
The policy gradient theorem states:
$\nabla_{\bm{\theta}}J(\bm{\theta})$ $=\mathbb{E}_{x,a}\!\left[\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(a|x)\,Q(x,a)\right].$ (43)
For the Bernoulli parameterization and noting that the policy function resembles sigmoid activation function the derivative of the log–policy can be simplified as
$\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(a|x)=\begin{cases}\displaystyle-\lambda_{p}(1-\pi_{0})&\text{if }x=0,a=0\\[6.0pt] \displaystyle\lambda_{p}\pi_{0}&\text{if }x=0,a=1\\[6.0pt] \displaystyle-(1-\pi_{1})&\text{if }x=1,a=0\\[6.0pt] \displaystyle\pi_{1}&\text{if }x=1,a=1\end{cases}$ (44)
where $\pi_{0}=\pi_{\bm{\theta}}(a=0\mid x=0)$ and $\pi_{1}=\pi_{\bm{\theta}}(a=0\mid x=1)$. Using the rewards in the environment, one can derive the analytic policy gradient vector
$\nabla_{\bm{\theta}}J(\bm{\theta})=\begin{bmatrix}-\lambda_{p}\,\pi_{0}(1-\pi_{0})(1-2\pi_{1})\\[6.0pt] -2(1-\pi_{0})\pi_{1}(1-\pi_{1})\end{bmatrix}$. (45)
The geometry of the policy manifold is captured by the FIM and as the states are independent of each other we can write the per state FIM as,
$F_{ii}=\mathbb{E}_{a\sim\pi(\cdot|x=i)}\big[\nabla_{\theta_{i}}\log\pi_{\theta}(a|x=i)\,\nabla_{\theta_{i}}\log\pi_{\theta}(a|x=i)^{\!\top}\big]$ (46)
So For State (x = 0): The Score functions for $a\in\{0,1\}$ can be expressed as:
$\frac{\partial}{\partial\theta_{0}}\log\pi_{\theta}(a=0\mid x=0)=-\lambda_{p}(1-\pi_{0}),\qquad\frac{\partial}{\partial\theta_{0}}\log\pi_{\theta}(a=1\mid x=0)=\lambda_{p}\pi_{0}$ (47)
Hence,
$F_{00}$ $=\pi_{0}[-\lambda_{p}(1-\pi_{0})]^{2}+(1-\pi_{0})[\lambda_{p}\pi_{0}]^{2}$ (48)
[…]
Similarly for state (x = 1): The Score functions for $a\in\{0,1\}$ ca be expressed as:
$\frac{\partial}{\partial\theta_{1}}\log\pi_{\theta}(a=0\mid x=1)=-(1-\pi_{1}),\qquad\frac{\partial}{\partial\theta_{1}}\log\pi_{\theta}(a=1\mid x=1)=\pi_{1}$ (52)
Hence,
$F_{11}$ $=\pi_{1}[-(1-\pi_{1})]^{2}+(1-\pi_{1})[\pi_{1}]^{2}$ (53)
[…]
Owing to independence between states, $F(\bm{\theta})$ is diagonal:
$F(\bm{\theta})=\begin{bmatrix}\lambda_{p}^{2}\,\pi_{0}(1-\pi_{0})&0\\[4.0pt] 0&\pi_{1}(1-\pi_{1})\end{bmatrix}$. (57)
Then the NPG update direction is obtained by preconditioning the gradient with the inverse Fisher matrix,
$\dot{\bm{\theta}}_{\text{NPG}}=F(\bm{\theta})^{-1}\,\nabla_{\bm{\theta}}J(\bm{\theta})=\begin{bmatrix}\dfrac{-(1-2\pi_{1})}{\lambda_{p}}\\[8.0pt] -2(1-\pi_{0})\end{bmatrix}$. (58)
To reduce computation while preserving approximate invariance, the SM approach uses a low–rank inverse metric $\tilde{F}^{-1}$ defined in terms of damping factor $\lambda$ and exact FIM $F(\theta)$ capturing dominant curvature direction,
$\tilde{F}^{-1}$ $=\frac{1}{\lambda}\left({\bm{I}}-\frac{F(\bm{\theta})}{\lambda+Tr(F(\bm{\theta}))}\right)$ (59)
Hence the resulting SM update direction can be written as:
$\dot{\bm{\theta}}_{\text{SM}}$ $=\tilde{F}^{-1}\,\nabla_{\bm{\theta}}J(\bm{\theta})$ (61)
$=\frac{1}{\lambda}\left({\bm{I}}-\frac{F(\bm{\theta})}{\lambda+Tr(F(\bm{\theta}))}\right)\nabla_{\bm{\theta}}J(\bm{\theta})$ (62)
$=\frac{1}{\lambda}\left(\nabla_{\bm{\theta}}J(\bm{\theta})-\frac{F(\bm{\theta})\nabla_{\bm{\theta}}J(\bm{\theta})}{\lambda+Tr(F(\bm{\theta}))}\right)$ (63)
$=\begin{bmatrix}\dfrac{1}{\lambda}\!\left(-\lambda_{p}\,\pi_{0}(1-\pi_{0})(1-2\pi_{1})\right)-\dfrac{\lambda_{p}^{3}\,\pi_{0}^{2}(1-\pi_{0})^{2}(1-2\pi_{1})}{\lambda^{2}+\lambda\!\left[\lambda_{p}^{2}\,\pi_{0}(1-\pi_{0})+\pi_{1}(1-\pi_{1})\right]}\\ \dfrac{1}{\lambda}\!\left(-2(1-\pi_{0})\pi_{1}(1-\pi_{1})\right)-\dfrac{2\,\pi_{1}^{2}(1-\pi_{1})^{2}(1-\pi_{0})}{\lambda^{2}+\lambda\!\left[\lambda_{p}^{2}\,\pi_{0}(1-\pi_{0})+\pi_{1}(1-\pi_{1})\right]}\end{bmatrix}$ (64)
A.6 Additional Results
Figure 8: Raw results on three classic control tasks. These correspond to the results reported in Fig. 6, but without binning or smoothing.
Figure 9: Results on six MuJoCo tasks. These correspond to the results reported in Fig. 7, but without binning or smoothing.
Figure 10: Average log-probabilities of the actions taken on three classic control tasks. All results are averaged over $5$ random seeds, with shaded areas showing standard deviation.
Figure 11: Average log-probabilities of the actions taken on six MuJoCo tasks. All results are averaged over $5$ random seeds, with shaded areas showing standard deviation.
A.7 Hyper-parameters
We consider the following hyper-parameters:
  • $\eta$ (actor) : step size for the policy‑network update.
  • $\alpha$ (critic) : learning rate for the value network, optimized with Adam.